Maxine:

Andrew invested $20,000 in mutual funds and received a sum of $35,000 at the end of the investment period. Calculate the ROI. A. 75% B. 65% C. 55%

XioGonz:

1 attachment
smackz:

ROI= (gain from investment - cost of investment)/.cost of investment = (35,000-20,000)/20,000 = 15,000/20,000 = ? ^ what is that in a decimal to precentage form cuz that part is ROI